A car waiting on a signal line starts moving with a uniform acceleration of 4m/s^2 , when traffic lights turn green . at the same time a motor cyclist moving with a constant speed of 36 km/h passes the car.find out
a)the distance from signal line at which the car driver over takes the motorcyclist and,
b) speed of car at that time.

Answer by Alan3354(69443)

A car waiting on a signal line starts moving with a uniform acceleration of 4m/s^2 , when traffic lights turn green . at the same time a motor cyclist moving with a constant speed of 36 km/h passes the car.find out
a)the distance from signal line at which the car driver over takes the motorcyclist and,
b) speed of car at that time.
------
36 km/hr = 10 m/sec
-------
find out
a)the distance from signal line at which the car driver over takes the motorcyclist
s(t) for the car = at^2/2 = 2t^2 meters
s(t) for the MC = 10t meters
------
2t^2 = 10t
t = 0 (at the light)
t = 5 seconds.
s = 50 meters
======================================
b) speed of car at that time.
----------
v = at = 4*5
v = 20 m/sec